# Sparrow: Data-Efficient Video-LLM with Text-to-Image Augmentation

### v4

@@ Abstract
Recent years have witnessed the success of Multimodal Large Language Models (MLLMs) in the vision understanding domain. The success of these models can largely be attributed to the dominant scaling law, which states that larger parameter sizes and data volumes contribute to better performance. Notably, data scaling has mainly been powered by automatic data pipelines, which center around the self-instruction of LLMs. The paradigm has been taken for granted for quite some time, but the study of the effectiveness of scaling with these data has been neglected for a long time. In this context, this work revisits scaling with synthetic data and focuses on developing video-LLMs from a data-centric perspective. Our main study approach is fine-tuning pre-trained image-LLMs with video data and investigating learning efficiency through data scaling. Results from our preliminary experiments reveal a low learning efficiency phenomenon when simply scaling up video data samples, which, through our probing, can be ascribed to a lack of instruction diversity. Aiming at this issue, we propose a data augmentation method called Sparrow, which synthesizes video-like samples from pure text instruction data. Mixing these synthetic samples with the video data enables a more efficient training scheme. Through comprehensive experiments, we demonstrate that our proposed method achieves performance comparable to or even superior to baselines trained with many more samples. Meanwhile, we find that incorporating these synthetic samples can boost the performance of long video understanding without training with long video data. The code and data examples are available at this link.
@@ 1 Introduction
The past few years have seen the rapid progress of Multimodal Large Language Models (MLLMs) [2, 3, 4]. Apart from solving traditional vision tasks (such as VQA), these models also excel in following user instructions and generalizing to new tasks. A mainstream paradigm for developing such models takes a two-stage training strategy. The first stage, pretraining, mainly serves to align vision modality with text and inject various kinds of visual knowledge into the model. In this stage, large-scale datasets of text-image pairs are often used, such as LAION [5] and CC [6], comprising a large proportion of the total compute and injecting abundant vision knowledge into models. Some methods also incorporate OCR and detection-related data to improve foundational capabilities [7, 4]. The second stage, instruction fine-tuning, adapts models to accommodate various tasks and helps generalize to new instructions. Training in this stage typically involves instruction data obtained from self-instruction or adaptation of task-specific datasets (e.g. VQA and chart understanding datasets). Recently, researchers have shifted their focus from single-image models to more advanced ones that support video understanding. Borrowing successful experience from developing image models, some video counterparts are typically trained from scratch, following a similar two-stage training paradigm [8, 9]. Apart from this path, some researchers utilize pre-trained image-LLMs instead. Typical approaches include zero-shot inference [10, 11, 12] and further fine-tuning [13, 9, 14, 8].
@@
Notably, the success of these models can be largely ascribed to the formidable scaling law, which puts emphasis on scaling up parameter size or data volume for better model performance. For the data aspect, the scaling has mainly been driven by automatic data engines, which synthesize massive amounts of data without human labor. Nevertheless, the characteristics of learning from these synthesized video data stand out as a critical yet underexplored topic. Thus, in this work, we investigate the learning characteristics of video-LLM more deeply from a data scaling perspective. Our preliminary data scaling experiments reveal a low data efficiency problem, that is, the performance gains from utilizing multiple times more data is marginal. An inspection of data characteristics suggests this might be due to a lack of instruction diversity in the training corpus. To address this issue, we propose a data augmentation method, dubbed Sparrow [^1] Inspiration taken from the swiftness of sparrows., to enrich the diversity of instruction. The basic idea is to synthesize video-like samples from textual data and mix these synthetic data with the video samples. Specifically, we use existing text instruction data whose sample comprises a (long-context, instruction, answer) triplet. The long-context part is split into multiple segments and then further transformed into images, while the instruction and answer stay intact. Processed in this way, the synthetic samples have the same structure as video instruction data and can be incorporated seamlessly into the training corpus.
@@ 2.1 Multimodal Large Language Models
Image-LLMs. To develop image-LLMs, the mainstream approach is to build upon powerful pre-trained LLMs and extend LLMs with the capability to perceive and reason with images [3, 4]. Based on a two-stage training recipe, i.e. image-text alignment training and instruction tuning, the developed model can fulfill a wide range of multimodal user queries and present its answers in user-friendly natural language sentences.
Video-LLMs. Following the success of image-LLMs, subsequent endeavors aim to expand the triumph to more intricate video understanding. Works like Video-ChatGPT [15], VTimeLLM [16], PLLaVA [17] and LLaVA-NeXT-Video [18] attempt further fine-tune image-LLMs to enhance video understanding capability. Other research [13, 9, 14, 8] explores training from pre-trained LLM, following the basic alignment-then-finetuning paradigm similar to image-LLM. These approaches usually involve joint training that mixes image and video data in the training corpus. In this study, we build upon pre-trained image-LLMs and enhance video understanding capabilities through fine-tuning.
@@ 2.2 Evaluation of Video Understanding
Early methods [10, 15] are generally evaluated on more traditional benchmarks like MSVD-QA [19], TGIF-QA [20] and ActivityNet-QA [21]). These benchmarks are generally domain-specific and focus on certain basic skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage (especially in longer videos) and skill coverage. Moreover, the questions asked often involve shallow perception without deeper reasoning.
Recently, with the rise of benchmarks specifically designed for MLLMs [22, 9, 23, 24], a more in-depth and comprehensive evaluation has become more accessible. Compared to previous traditional benchmarks, these newly developed benchmarks are generally more challenging, often entailing composite skills and a finer-grained understanding of the video (e.g. the plot in the movie or causal relationships between events), and can be much longer in duration (e.g. up to 60 minutes in the Video-MME benchmark). In this work, our study adopts these newly developed video benchmarks.
@@ 3.1.3 Training Datasets
ShareGemini-Webvid-core100k. It is a video caption dataset with 100K samples in total. The videos are curated from WebVid [33], a web-scale video-caption dataset covering open domains and general subjects. Regarding duration, the dataset mainly contains short videos with lengths shorter than 30 seconds.
@@ 3.2 Evaluation Setup
To evaluate the model capabilities in an efficient and comprehensive way, we use Video-MME [22], MVBench [9], and TempCompass [23] as our benchmarks. We do not use traditional video-QA benchmarks (e.g. MSVD-QA [19], TGIF-QA [20], ActivityNet-QA [21]) since these benchmarks are generally limited to a small coverage of domains, task types, and video lengths. Moreover, the questions asked often involve shallow perception without deeper reasoning since early models generally lack reasoning capacity, whereas recent LLM-based models excel. We illustrate more about the benchmarks used as follows:
@@
TempCompass focuses on fine-grained temporal aspects, such as action, speed, and attribute change. The videos and meta-information are manually collected, after which annotations are generated by LLMs with the aid of meta-information. We use the multiple-choice QA (MCQ) format to align with other benchmarks.
To ensure robust and efficient judging of model answers, we use a combination of exact matching and LLM matching for assessment. More details about the implementation of this evaluation scheme are available in Appendix A.1.
@@ 3.3.1 Low Learning Efficiency Issue
Figure 3: Scaling performance of fine-tuning InternVL with different data volumes and types on different benchmarks. We consider fine-tuning with ShareGemini, Video-ChatGPT, and a mix of both with a 1:$1:11$ : 1 sampling ratio. Training samples are measured in K, where 0 sample indicates zero-shot inference.
Figure 4: t-SNE plot of instruction distribution of video datasets - ShareGemini and Video-ChatGPT. 5,000 instructions are sampled from each dataset. The instruction diversity of these two datasets is relatively limited, leading to low learning efficiency.
@@ 4.1 Design Concept
Figure 5: A conceptual illustration of our scheme for synthesizing video-like samples from text data. For each (long-context, instruction, answer) triplet, we split the long context into segments and transcribe these segments into a series of images, simulating the structure of a video. Zoom in for better visibility.
Since currently available video data can be limited in instruction diversity, and annotating high-quality video data is costly, we aim to expand the instruction diversity by incorporating new synthetic data. A rich source of instruction data lies in the text domain, and they can effectively complement the vision domain. Nevertheless, there is inherently a modality gap between the text and visual domains. To better utilize these data, we bridge the modality gap by synthesizing images with the text. Fig. 5 illustrates our overall workflow for synthesis of data.
Figure 6: t-SNE plot of instruction distribution after applying our proposed methods. By incorporating text data, i.e. LongAlpaca [41] and LongQLoRA [41], the overall distribution becomes more diverse.
Our proposed scheme enjoys three benefits: (1) Mixing in text data can effectively enrich the instruction diversity (Fig. 6), thus improving the learning efficiency for video fine-tuning; (2) Images synthesized from text can emulate the 1D temporal structure of video frames since text segments are generally correlated in the context, thus mitigating the gap between common video samples and synthetic ones; (3) Text data are easier to collect than video samples. Thus, utilizing synthetic data can be economical.
@@ 4.2 Implementation Details
Each text sample is a (long-context, question, answer) triplet. For example, the long context can be a section of a book or an academic paper, while the instruction and the answer are centered around the context, e.g. an inquiry to give a synopsis, or questions related to the paper. After the data transformation process, each sample is a video-like (images, question, answer) triplet, where long-context information is transformed into a series of images, and the question and answer stay unchanged.
The key to the data synthesis procedure is synthesizing images with pure text. Specifically, for each (long-context, question, answer) triplet, we divide the context information into multiple segments according to word counts (set to 115 empirically) using an open-sourced NLP toolkit [^2] https://www.nltk.org/. These text chunks are then transformed into a sequence of images separately. Specifically, each text chunk is embedded into a blank image with a white background. This is achieved using bitmap font with the ImageFont module of Pillow library [^3] https://pillow.readthedocs.io/en/stable/. Each image is 448x448 pixels in size, and the font is 20 pt large, black color, Arial Regular type. We use a bounding box to control the layout, leaving a margin of 20 pixels on each side, so each line of the text has roughly the same width. After the transformation, the structure of these synthesized data is exactly the same as that of video samples, and we can seamlessly incorporate the synthetic data into the video training corpus.
@@ 5 Evaluation on Proposed Methods
This section includes experimental results and discussions of our proposed method, including (1) a comparison with mainstream methods, (2) an ablation study on data mixes, and (3) an examination of key properties, including data scaling performance and gains in long video understanding.
@@ 5.1 Comparison with Mainstream Methods
We compare the results with some representative proprietary models, including GPT-4V [42], Claude 3.5 Sonnet [43], GPT-4o [44], Gemini 1.5 Pro [45], and open-source video-LLMs of similar LLM parameter size, including Video-LLaVA [13], VideoChat2 [9], Chat-UniVi-v1.5 [14], VideoLLaMA 2 [8], VITA [46], VITA-1.5 [48], and Kangaroo [47], as summarized in Tab. 1.

### v5

@@ Abstract
Recent years have seen the success of Multimodal Large Language Models (MLLMs) in the domain of vision understanding. The success of these models can largely be attributed to the dominant scaling law, which states that larger parameter sizes and data volumes contribute to better performance. Notably, data scaling has been primarily driven by automatic data pipelines, which focus on the self-instruction of LLMs. The paradigm has been taken for granted for quite some time, but the study of the effectiveness of scaling with these data has been neglected for a long time. In this context, this work revisits scaling with synthetic data and focuses on developing video-LLMs from a data-centric perspective. Our primary study approach involves fine-tuning pre-trained image-LLMs with video data and examining learning efficiency through data scaling. Results from our preliminary experiments reveal a low learning efficiency phenomenon when simply scaling up video data samples, which, through our probing, can be ascribed to a lack of instruction diversity. Aiming at this issue, we propose a data augmentation method called Sparrow, which synthesizes video-like samples from pure text instruction data. Mixing these synthetic samples with the video data enables a more efficient training scheme. Through comprehensive experiments, we demonstrate that our proposed method achieves performance comparable to or even superior to that of baselines trained with significantly more samples. Meanwhile, we find that incorporating these synthetic samples can enhance the performance of long video understanding without requiring training on long video data. The code and data examples are available here.
@@ 1 Introduction
The past few years have seen the rapid progress of Multimodal Large Language Models (MLLMs) [2, 3, 4, 5, 6, 7, 8, 9]. Apart from solving traditional vision tasks (such as VQA), these models also excel in following user instructions and generalizing to new tasks. A mainstream paradigm for developing such models takes a two-stage training strategy. The first stage, pretraining, mainly serves to align vision modality with text and inject various kinds of visual knowledge into the model. In this stage, large-scale datasets of text-image pairs are often used, such as LAION [10] and CC [11], comprising a large proportion of the total compute and injecting abundant vision knowledge into models. Some methods also incorporate OCR and detection-related data to improve foundational capabilities [12, 13]. The second stage, instruction fine-tuning, adapts models to accommodate various tasks and helps generalize to new instructions. Training in this stage typically involves instruction data obtained from self-instruction or adaptation of task-specific datasets (e.g., VQA and chart understanding datasets). Recently, researchers have shifted their focus from single-image models to more advanced ones that support video understanding. Borrowing successful experience from developing image models, some video counterparts are typically trained from scratch, following a similar two-stage training paradigm [14, 15]. Apart from this path, some researchers utilize pre-trained image-LLMs instead. Typical approaches include zero-shot inference [16, 17, 18] and further fine-tuning [19, 15, 20, 14].
@@
Notably, the success of these models can be largely ascribed to the formidable scaling law, which puts emphasis on scaling up parameter size or data volume for better model performance. For the data aspect, the scaling has mainly been driven by automatic data engines, which synthesize massive amounts of data without human labor. Nevertheless, the characteristics of learning from these synthesized video data stand out as a critical yet underexplored topic. Thus, in this work, we investigate the learning characteristics of video-LLM more deeply from a data scaling perspective. Our preliminary data scaling experiments reveal a low data efficiency problem, that is, the performance gains from utilizing multiple times more data are marginal. An inspection of data characteristics suggests this might be due to a lack of instruction diversity in the training corpus. To address this issue, we propose a data augmentation method, dubbed Sparrow [^1] Inspiration taken from the swiftness of sparrows., to enrich the diversity of instruction. The basic idea is to synthesize video-like samples from textual data and mix these synthetic samples with the video samples. Specifically, we use existing text instruction data whose sample comprises a (long-context, instruction, answer) triplet. The long-context part is split into multiple segments and then further transformed into images, while the instruction and answer stay intact. Processed in this way, the synthetic samples have the same structure as video instruction data and can be incorporated seamlessly.
@@ 2.1 Multimodal Large Language Models
Image-LLMs. To develop image-LLMs, the mainstream approach is to build upon powerful pre-trained LLMs and extend LLMs with the capability to perceive and reason with images [21, 13]. Based on a two-stage training recipe, i.e., image-text alignment training and instruction tuning, the developed model can fulfill a wide range of multimodal user queries and present its answers in user-friendly natural language sentences.
Video-LLMs. Following the success of image-LLMs, subsequent endeavors aim to expand the triumph to more intricate video understanding. Works like Video-ChatGPT [22], VTimeLLM [23], PLLaVA [24], and LLaVA-NeXT-Video [25] attempt to further fine-tune image-LLMs to enhance video understanding capability. Other research [19, 15, 20, 14] explores training from pre-trained LLM, following the basic alignment-then-finetuning paradigm similar to image-LLM. These approaches usually involve joint training that mixes image and video data in the training corpus. In this study, we build upon pre-trained image-LLMs and enhance video understanding capabilities through fine-tuning.
2.2 Long Video Understanding
A fundamental challenge in long video comprehension lies in the effective modeling of long video frame sequences. To tackle this problem, two major technical directions have been actively investigated, namely context window extension and efficient video modeling.
Context Window Extension. To accommodate a longer sequence of video frames with completeness, an intuitive approach is to extend the context window of the LLM backbone. Previous works have mainly adopted continued pretraining on the LLM before video training. For instance, LongVA [26], LongVILA [27], and Kangaroo [28] perform continued training on long text corpora to accommodate more video tokens.
Efficient Video Modeling. With the observation that visual information is often redundant in spatial and temporal dimensions, and that key frames are sparse in videos, another line of work explores efficient modeling of long videos. These works try to compress the video token sequences to fit in the original context window of language models. Specifically, LLaMA-VID [29] proposes an attention-based module to compress each video frame into two tokens. MovieChat [30] designs a memory mechanism to cache video context and combines both short-term and long-term memory. TimeChat [31] combines image Q-Former and video Q-Former to compress video tokens within single frames and sliding windows of video, respectively. Other works seek to retrieve frames most relevant to the user prompt and thus directly cut down the input length. For example, KeyVideoLLM [32] utilizes a pretrained CLIP encoder to calculate similarities between the query and video frames, and selects the most relevant frames based on the similarities. Similarly, Video-RAG [33] further incorporates external tools to extract auxiliary texts from the video and augments the input with more multimodal context, such as OCR, ASR, and object detection information.
@@ 2.3 Evaluation of Video Understanding
Early methods [16, 22, 34] are generally evaluated on traditional benchmarks like MSVD-QA [35], TGIF-QA [36] and ActivityNet-QA [37]). These benchmarks are generally domain-specific and focus on certain basic skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage (especially in longer videos) and skill coverage. Moreover, the questions asked often involve shallow perception without deeper reasoning.
Recently, with the rise of benchmarks specifically designed for MLLMs [38, 15, 39, 40], a more in-depth and comprehensive evaluation has become more accessible. Compared to previous traditional benchmarks, these newly developed benchmarks are generally more challenging, often entailing composite skills and a finer-grained understanding of the video (e.g., the plot in the movie or causal relationships between events), and can be much longer in duration (e.g., up to 60 minutes in the Video-MME benchmark). In this work, our study adopts these newly developed video benchmarks.
@@ 2.4 Textual Data for Video Understanding
3 Problem Formulation
We focus on mainstream MLLMs architectures [2], which typically consist of a vision encoder, a projector, and an LLM backbone. Given a video $\mathbf{V}$ downsampled to $T$ frames as $\mathbf{F}=\{{f_{i}}\}_{i=1}^{T}$, frame-level features are extracted via $\mathbf{E}=\{\mathrm{E}_{i}\}_{i=1}^{T}=\texttt{ViT}(\mathbf{F})$, where ViT denotes the vision encoder (e.g., a pre-trained model such as CLIP [43]). Each $\mathrm{E}_{i}\in\mathbb{R}^{(H\times W)\times C}$ represents the visual tokens of the $i$-th frame, with $H$, $W$ being the spatial dimensions and $C$ the feature dimension. These visual features are then projected into the text embedding space via a projector module, typically an MLP, yielding $\mathbf{\hat{E}}=\texttt{Proj}(\mathbf{E})$. The transformed vision features are subsequently concatenated with the text embeddings of a user query $\mathbf{Q}$ to form a joint token sequence $[\textrm{w}_{V};\textrm{w}_{T}]$. This combined sequence is fed into the LLM backbone, which generates a natural language response auto-regressively:
$p(w_{o}|\textrm{w}_{V},\textrm{w}_{T})\sim\prod_{t=1}^{L}P(w_{t}|w_{<t},\textrm{w}_{V},\textrm{w}_{T}),$ (1)
where $\textrm{w}_{o}=\{w_{o,t}\}_{t=1}^{L}$ denotes the output token sequence of length $L$. Here, $\textrm{w}_{V}$ corresponds to the vision tokens processed through the encoder and projector, while $\textrm{w}_{T}$ refers to the tokenized representation of the user query.
In terms of training data composition, each instance is structured as a triplet $(\mathbf{V},\mathbf{Q},\mathbf{A})$, where $\mathbf{Q}$ denotes a natural language instruction and $\mathbf{A}$ is the corresponding textual answer. Let $\mathcal{D}$ represent the data distribution over such triplets. The model is trained to generate the answer sequence $\mathbf{A}$ conditioned on both the video input $\mathbf{V}$ and the instruction $\mathbf{Q}$. Formally, we aim to minimize the expected negative log-likelihood loss over the model answer:
$\mathcal{L}(\theta)=\mathbb{E}_{(\mathbf{V},\mathbf{Q},\mathbf{A})\sim\mathcal{D}}\left[-\log\,p_{\theta}(\mathbf{A}\mid\mathbf{V},\mathbf{Q})\right].$ (2)
where $\theta$ denotes the model parameters and $p_{\theta}(\cdot)$ is the conditional output distribution of the MLLM.
@@ 4.1.3 Training Datasets
ShareGemini-Webvid-core100k. This dataset comprises 100K video-caption pairs in total. The videos in the dataset are sourced from WebVid [48], a web-scale video-caption dataset spanning diverse open-domain topics. In terms of temporal duration, the majority of videos are short-form, with lengths under 30 seconds.
@@ 4.2 Evaluation Setup
To evaluate the model capabilities in an efficient and comprehensive way, we use Video-MME [38], MVBench [15], and TempCompass [39] as our benchmarks. We do not use traditional video-QA benchmarks (e.g., MSVD-QA [35], TGIF-QA [36], ActivityNet-QA [37]) since these benchmarks are generally limited to a small coverage of domains, task types, and video lengths. Moreover, the questions asked often involve shallow perception without deeper reasoning since early models generally lack reasoning capacity, whereas recent LLM-based models excel. We illustrate more about the benchmarks used as follows:
@@
TempCompass is designed to evaluate models’ temporal understanding capabilities, encompassing temporal aspects such as action, speed, and attribute change. The videos and corresponding meta-information are manually collected, followed by LLM-generated annotations. We use the multiple-choice QA (MCQ) format to align with other benchmarks.
To ensure robust and efficient judging of model answers, we use a combination of exact matching and LLM matching for assessment. More details about the implementation of this evaluation scheme are available in Appendix A.
@@ 4.3.1 Low Learning Efficiency Issue
Figure 3: Scaling performance of fine-tuning InternVL with different data volumes and types on different benchmarks. We consider fine-tuning with ShareGemini, Video-ChatGPT, and a mix of both with a $1:1$ sampling ratio. Training samples are measured in K, where 0 sample indicates zero-shot inference.
Figure 4: t-SNE plot of instruction distributions for video datasets – ShareGemini and Video-ChatGPT. We sample 5,000 instructions from each dataset for visualization. The relatively limited instruction diversity observed in both datasets may hinder learning efficiency during fine-tuning.
@@ 5.1 Design Concept
Figure 5: Conceptual illustration of our proposed scheme. (a) We illustrate the structures of text, synthetic, and video samples. Synthetic video-like samples are generated from textual data, with a structure mimicking real video samples. Specifically, for each (long context, instruction, answer) triplet, the long context part is split into segments, which are then transcribed into a sequence of text-rich images, simulating a video. (b) The domain differences between real and synthetic sources contribute to greater instruction diversity when used jointly for training. (c) The sequential nature of synthetic image sequences aligns well with the structure of real video data, thus enabling a unified training format.
Since currently available video data can be limited in instruction diversity, and annotating high-quality video data is costly, we aim to expand the instruction diversity by incorporating new synthetic data. A rich source of instruction data lies in the text domain, and it can effectively complement the vision domain. Nevertheless, there is inherently a modality gap between the text and visual domains. To better utilize these data, we bridge the modality gap by synthesizing images with the text. Fig. 5 illustrates the overall data synthesis pipeline and characteristics of our scheme. Our proposed scheme enjoys three benefits: (1) Mixing in text data can effectively enrich the instruction diversity (Fig. 6), thus improving the learning efficiency for video fine-tuning; (2) Images synthesized from text can emulate the 1D temporal structure of video frames since text segments are generally correlated in the context, thus mitigating the gap between common video samples and synthetic ones; (3) Text data is easier to collect than video samples. Thus, utilizing synthetic data can be economical.
Figure 6: t-SNE plot of instruction distribution after applying our proposed method.
@@ 5.2 Implementation Details
Each text sample is a (long-context, question, answer) triplet. For example, the long context can be a section of a book or an academic paper, while the instruction and the answer are centered around the context, e.g., an inquiry to give a synopsis, or questions related to the paper. After the data transformation process, each sample is a video-like (images, question, answer) triplet, where long-context information is transformed into a series of images, and the question and answer stay unchanged.
The key to the data synthesis procedure is synthesizing images with pure text. Specifically, for each (long-context, question, answer) triplet, we divide the context information into multiple segments according to word counts (set to 115 empirically) using an open-sourced NLP toolkit [^2] https://www.nltk.org/. These text chunks are then transformed into a sequence of images separately. Specifically, each text chunk is embedded into a blank image with a white background. This is achieved using a bitmap font with the ImageFont module of the Pillow library [^3] https://pillow.readthedocs.io/en/stable/. Each image is 448x448 pixels in size, and the font is 20 pt large, black color, Arial Regular type. We use a bounding box to control the layout, leaving a margin of 20 pixels on each side, so each line of the text has roughly the same width. Following the transformation, the synthetic data aligns structurally with the video samples, enabling its direct and seamless incorporation into the video training sets.
@@ 6 Evaluation on Proposed Methods
This section includes experimental results and discussions of our proposed method, including (1) a comparison with mainstream methods, (2) an ablation study on data mixes, (3) an examination of key properties, including data scaling performance and gains in long video understanding, and (4) an in-depth analysis of model performance, including qualitative results as well as performance breakdown across various task types.
@@ 6.1 Comparison with Mainstream Methods
We compare the our method with some representative proprietary models, including GPT-4V [56], Claude 3.5 Sonnet [57], GPT-4o [58], Gemini 1.5 Pro [59], and open-source video-LLMs of similar LLM parameter size, including Video-LLaVA [19], VideoChat2 [15], Chat-UniVi-v1.5 [20], VideoLLaMA 2 [14], VITA [60], VITA-1.5 [61], and Kangaroo [28]. The results are summarized in Table 1.